# Current Trends in Digital Twin Development, Maintenance, and Operation: An Interview Study

### v3

@@ 1 Introduction
Digital Twins (DTs) have captured the interest of industry and academia in recent years, because of their promise to help to better understand, monitor and improve systems. The concept of DTs often encompasses the notion of a real-world entity and a digital counterpart that mimics certain aspects of the former. We believe that both industry and academia are playing a crucial role in further developing the DT concept as well as its design, development, operation, and maintenance. Although growing in popularity, the concepts and practices used around DTs show significantly differences. According to Zhang et al. [31] there is no general consensus on the nature of the real-world entity, the required fidelity of the digital counterpart, or the terms used to refer to these entities.
@@
In recent times, an increasing number of domains are developing and/or using DTs to address use-cases spanning over a large spectrum of complexity. A significant number of these DTs are related to complex industrial systems that utilize cutting-edge technologies to execute complex tasks with an extremely high degree of precision and efficiency. Various cross-domain and heterogeneous models plays a central role within these systems and the use of model-driven engineering (MDE) techniques are paramount in the design, development, and maintenance of these models, which is clearly apparent from the current literature [3, 29, 11, 14]. We believe that the role of models and MDE is equally crucial in the development of the digital counterpart or virtual entity (VE) of these physical industrial systems in the context of DTs. Our viewpoint is based on three facts: firstly, models are widely recognised as abstractions of real-world concepts, which aligns with the overarching objective of VEs and, therefore, integration of diverse cross-domain models into a VE is logical; secondly, reuse of existing models and related artifacts can significantly improve the development speed of the corresponding VE; finally, the general benefit of MDE, which makes it increasingly popular [11], also apply to the development of VEs.
@@
Figure 1: Depiction of the concept of digital twin as introduced by Grieves [12] where he conceptualized a virtual space or virtual entity (VE) and synchronisation with its physical counterpart. Our research focuses on the VE.
@@
As model-based systems, VEs consist of cross-domain and heterogeneous models, which must be (KC1) kept consistent, (KC2) properly orchestrated, and (KC3) validated for correct behavior. Although these activities are identified as key challenges in [28], they have not received sufficient attention in the current literature, as far as we are aware. We believe this shortfall can be attributed to the vast diversity of DTs and their rapid large-scale adoption both in industry and academia. In this research, we aimed to better understand the concepts, practices, and methods around the design, development, and maintenance of VEs. To achieve this, we interviewed 19 individuals from industry and academia who are all involved in the design, development, maintenance, and use of DTs of complex industrial systems. The interviews and subsequent analysis provided us with practical insights into diverse understanding of DTs and practices around the three earlier mentioned challenges. Our work was guided by the following seven research questions (RQ):
@@ 2 Background
The concept of Digital Twins (DTs) was introduced by Grieves [12] in 2003; Grieves modelled DTs with three dimensions i.e., the physical entity, virtual model and connection, which facilitates the physical–virtual interaction. Since then researchers and practitioners have used DT as an umbrella term to refer to something from a simple simulation to a complex virtual entity closely mimicking a real-world counterpart. For example, Bielefeldt et al. [13] focus on ultra-realistic multi-physical computational models in their definition of DT, whereas El Saddik [18] defines a DT as a digital replica of a physical entity whether living or non-living.
Tao et al. [19] extended the original DT model by Grieves and proposed a five dimensional (5-D) model depicted in Figure 2, i.e., $M_{DT}$ = (PE, VE, Ss, DD, CN) where PE refers to the physical (actual) real-world entity with various functional subsystems, VE is the corresponding high-fidelity digital model that reproduces certain abilities and properties of the PE, Ss represents the services for PE and VE, DD encapsulates the domain knowledge–data from both PE & VE and their fusion, and finally, CN is the connection among parts of the DT.
@@
Figure 2: Five dimensional (5D) model of a DT proposed by Tao et al. [19]
Since the inception of DTs, research has focused on understanding the concept, the development of DT’s applications, or exploration of different implementation technologies used. An example is the work of Liu et al. [30] that analyses the concept of a DT, the technologies used, and DTs’ main industrial applications. That research is based on systematic literature review (SLR) analysing over 200 publications.
Tao et al. [22] analyses the state-of-the-art in development and applications of DTs, aiming to outline key technologies enabling DT development, classify current and future applications, and lay out possible gaps and challenges. The research used an SLR, analyzing 50 papers and eight patents.
Sharma et al. [40], based on an SLR of over 80 papers, analysed and proposed solutions for the research and implementation gaps of DT technology, such as IoT (internet of things), machine learning and data. This research concluded that regulation and security mechanisms are essential for the proper implementation of DTs due to its cross-domain nature. They also concluded that there are multiple technical and domain specific challenges that require more research to be resolved.
Gürdür et al. [44] explores how DTs can help the infrastructure industry. The research methodology used semi-structured expert interviews with non-technical executives from industry in the UK. This approach allowed the researchers to collect their opinions, related to non-technical challenges, on the value of DTs.
Dalibor et al. [34] conducted an SLR on 356 papers, analysing DTs with a bottom-up approach exploring different implementations to investigate expected DT properties and how DTs are deployed, operated and evaluated. In addition, the authors developed a DT feature model. They explored different implementation techniques, tooling and development processes.
@@ 3 Motivation and contribution
Although Grieves [12] introduced the concept of DT more than two decades ago, practitioners are discovering its potential only recently, resulting into an increase of activities and practices around DTs. An increasing number of DTs are being developed and used in domains such as automotive, healthcare, manufacturing, and construction. The application of these DTs extends across a wide spectrum of use-cases, ranging from basic monitoring to highly intricate control systems and predictive maintenance applications. Due to the diversity of use-cases and domains involved, the tools, methods, and practices around the design, development, and maintenance of DTs vary significantly. The central objective of this research is to understand and study this diversity. Our work concentrates on the design, development, and maintenance of the virtual part of a DT, which we refer to as virtual entity (VE) earlier.
VE, which is often described as digital replication of certain capabilities of a physical entity, is essentially a complex software- and data-intensive system that encompasses various heterogeneous and multi-domain models [26, 22]. To maintain the continued operation of a DT, the corresponding models must be kept consistent with each other, they need to be orchestrated properly, and finally, the dynamic behaviour of both the individual models and the DT as a whole must be validated. Although these tasks are well-known and researched within the broader model-driven engineering (MDE) community, we are unaware of any work focusing on these topics in the context of DTs. This deficiency motivates our work to interview 19 DT practitioners from industry and academia, in order to acquire insight into current challenges, methods and practices related to the three previously mentioned activities (i.e., KC1–3) in the context of DTs.
@@ 4 Research methodology
Considering the exploratory nature of our research questions we opted for semi-structured interviews [24]. This provided sufficient flexibility for the participants to express themselves while allowing us to collect data on our topics of interest. In this section, we introduce and explain our research methodology, following Strandberg’s interview lifecycle [21], with the steps depicted in Figure 3. We expand on each step and explain the related activities in the following subsections.
@@ 4.1.2 Questionnaire design
We prepared a questionnaire to act as a guideline to keep the discussion in our semi-structured interviews focused. It was developed in line with the Interview Protocol Refinement (IPR) framework [15], comprising four phases:
@@
    Constructing an inquiry-based conversation: We categorized the interview questions into (1) background, (2) key, and (3) concluding ones. Based on this categorization and suggestions of Hove and Anda [4], we sorted them and rephrased some, enabling an inquiry-based conversation.
@@
Table 1 shows the resulting set of interview questions and their associations with the RQs. All the interview questions except for the first two are connected to one or more RQs. These two questions allowed us to start the conversation, get acquainted with the interviewee, and contributed to a conversation-like interview. Furthermore, with the final question we asked the interviewees’ opinion on Tao et al.’s 5-D DT model [19] (see Section 2). While asking this question, we showed an image of the 5-D model and briefly explained it. To avoid influencing the interviewees during the rest of the interview, we intentionally asked this question at the end.
@@
Research questions (RQs) normal-$\rightarrow$ 1 2 3 4 5 6 7 8
@@
1.
What is the domain of your company and what kind of services does it provide?
2.
How long have you been working and what is your current role?
@@
3.
What is your understanding of a DT?
$\times$
4.
How are you involved in the development or usage of the DT?
5.
What problems are you solving with your DT?
$\times$
6.
Is your DT for the entire process/system or a specific part?
$\times$
7.
What are the main parts of your DT? Could you shortly describe their role?
$\times$ $\times$
8.
Is there a physical counterpart of your DT? Does it communicate with the digital world? If yes, how?
$\times$ $\times$
9.
Did you build your DT from scratch or you reused some of the things that already existed? What kind of issues did you face if you reused something or by building from scratch?
$\times$ $\times$
10.
How is the data exchange between the models (and physical system) specified?
$\times$ $\times$ $\times$
11.
How is the sequence of execution of the models in your DT?
$\times$
12.
What issues, if any, did you face with the overall collective execution of models in your DT?
$\times$
13.
Which platforms/tools are you using to develop the digital twin?
$\times$ $\times$ $\times$ $\times$
14.
How often is the DT updated for bug fixes, improvement or similar reasons?
$\times$
15.
After each update did you face integration issues? If yes, what kind?
$\times$ $\times$
16.
How do you ensure various software elements can work together specially if multiple tools were used for development?
$\times$ $\times$ $\times$ $\times$ $\times$
17.
What properties/parameters did you validate to ensure an overall consistent behavior of your DT?
$\times$
18.
What tools and techniques did you use to validate these properties and parameters?
$\times$ $\times$
19.
What do you consider to be the general characteristics of your DT?
$\times$
@@
20.
How do you see the DT evolve in the future to solve additional problems?
$\times$
21.
What is your opinion on the 5-dimensional DT model from Tao et al. [19]?
$\times$
@@ 4.5 Data analysis
We further analysed the transcripts based on the thematic qualitative analysis methodology [23], the activities related to which are explained in this section. Each of these activities utilize outcomes from one or more of the previous activities. Figure 4 depicts the relationships among the activities and the related outcomes.
We used LaMa [43], a web-based tool for collaborative labeling and thematic analysis, to collaborate on this analysis. To restrict access and ensure data privacy, we deployed this platform locally.
@@ 4.5.2 Labeling of artifact and topic generation
After generating the artifacts for all transcripts, we labeled them using LaMa [43]. We define a label as a short text that sufficiently captures the core message of an artifact. To reduce bias in labeling, each artifact was labeled by two labelers. We resolved conflicting labels by agreeing on one label through discussion. During the labeling process, the labeler could use an existing label or create a new one. In LaMa, these labels were accompanied by a description explaining how and when a label should be used, which was crucial for reuse of existing labels. Moreover, while labeling we encountered artifacts that lacked sufficient information or context. We labeled these artifacts with two predefined labels: No value or Not understandable. An example is an artefact discussing an interviewee research policy in relation with their clients. This artefact was classified as No value, since it does not discuss any information related to DT development.
@@ 4.6 Replication package
The replication package for this interview study is available at [45]. Due to the agreement made with our interviewees, we must keep the interview transcripts confidential. Although this significantly constraints on the reproducibility and validation of the findings presented in this paper, we believe that the provided package can still significantly contribute to similar research endeavours. This package includes:
@@ 6.1.1 Virtual representation of an entity
Table 3: Correlation between DT components from interviewees and 5D model [19] addressed by industrial (#I) and academic (#A) interviewees.
@@ 6.1.2 Components of a DT
  • Purpose: Thirteen interviewees expressed that a DT should have a purpose and some mentioned that this purpose is the driving factor for DTs to be developed. On the contrary, one industrial interviewee explicitly mentioned that DTs should not have a specific purpose. This interviewee further discussed that DTs should not be developed with a specific purpose and they have a purposeless existence, which is in stark contrast to what the thirteen interviewees mentioned above. This interviewee clarified that once the DT is developed, it can be used for whichever purpose is needed. The purpose mentioned by the different interviewees can be correlated to the ‘services’ component of Tao et al. [19]’s 5D model of DTs (see Section 2).
@@ 6.1.3 Relation to the 5D DT model
As explained in Section 4.1.2, we collected and analyzed the opinions of our interviewees on the 5D DT model by Tao et al. [19] (explained in Section 2). This 5D DT model was selected in this research because it is a widely used model for DT representation and has been adopted for discussions on DTs by many. So, we wanted to understand the opinions of interviewees on this highly used and adopted model. In this section, we present our observations.
@@
Table 4: Agreement on 5D model of DT [19] by interviewees from academia (#A) and industry (#I).

### v4

@@ 1 Introduction
Digital Twins (DTs) have captured the interest of industry and academia in recent years, because of their promise to help to better understand, monitor and improve systems. The concept of DTs often encompasses the notion of a real-world entity and a digital counterpart that mimics certain aspects of the former. We believe that both industry and academia are playing a crucial role in further developing the DT concept as well as its design, development, operation, and maintenance. Although growing in popularity, the concepts and practices used around DTs show significantly differences. According to Zhang et al. [ZHANG2021151] there is no general consensus on the nature of the real-world entity, the required fidelity of the digital counterpart, or the terms used to refer to these entities.
@@
In recent times, an increasing number of domains are developing and/or using DTs to address use-cases spanning over a large spectrum of complexity. A significant number of these DTs are related to complex industrial systems that utilize cutting-edge technologies to execute complex tasks with an extremely high degree of precision and efficiency. Various cross-domain and heterogeneous models plays a central role within these systems and the use of model-driven engineering (MDE) techniques are paramount in the design, development, and maintenance of these models, which is clearly apparent from the current literature [baker2005, DEARAUJOSILVA2021101021, DAVIES201488, RODRIGUESDASILVA2015139]. We believe that the role of models and MDE is equally crucial in the development of the digital counterpart or virtual entity (VE) of these physical industrial systems in the context of DTs. Our viewpoint is based on three facts: firstly, models are widely recognised as abstractions of real-world concepts, which aligns with the overarching objective of VEs and, therefore, integration of diverse cross-domain models into a VE is logical; secondly, reuse of existing models and related artifacts can significantly improve the development speed of the corresponding VE; finally, the general benefit of MDE, which makes it increasingly popular [DAVIES201488], also apply to the development of VEs.
@@
Figure 1: Depiction of the concept of digital twin as introduced by Grieves [grieves2017digital] where he conceptualized a virtual space or virtual entity (VE) and synchronisation with its physical counterpart. Our research focuses on the VE.
@@
As model-based systems, VEs consist of cross-domain and heterogeneous models, which must be (KC1) kept consistent, (KC2) properly orchestrated, and (KC3) validated for correct behavior. Although these activities are identified as key challenges in [Models_meet_data], they have not received sufficient attention in the current literature, as far as we are aware. We believe this shortfall can be attributed to the vast diversity of DTs and their rapid large-scale adoption both in industry and academia. In this research, we aimed to better understand the concepts, practices, and methods around the design, development, and maintenance of VEs. To achieve this, we interviewed 19 individuals from industry and academia who are all involved in the design, development, maintenance, and use of DTs of complex industrial systems. The interviews and subsequent analysis provided us with practical insights into diverse understanding of DTs and practices around the three earlier mentioned challenges. Our work was guided by the following seven research questions (RQ):
@@ 2 Background
The concept of Digital Twins (DTs) was introduced by Grieves [grieves2017digital] in 2003; Grieves modelled DTs with three dimensions i.e., the physical entity, virtual model and connection, which facilitates the physical–virtual interaction. Since then researchers and practitioners have used DT as an umbrella term to refer to something from a simple simulation to a complex virtual entity closely mimicking a real-world counterpart. For example, Bielefeldt et al. [bielefeldt2015computationally] focus on ultra-realistic multi-physical computational models in their definition of DT, whereas El Saddik [el2018digital] defines a DT as a digital replica of a physical entity whether living or non-living.
Tao et al. [tao2017digital] extended the original DT model by Grieves and proposed a five dimensional (5-D) model depicted in Figure 2, i.e., $M_{DT}$ = (PE, VE, Ss, DD, CN) where PE refers to the physical (actual) real-world entity with various functional subsystems, VE is the corresponding high-fidelity digital model that reproduces certain abilities and properties of the PE, Ss represents the services for PE and VE, DD encapsulates the domain knowledge–data from both PE & VE and their fusion, and finally, CN is the connection among parts of the DT.
@@
Figure 2: Five dimensional (5D) model of a DT proposed by Tao et al. [tao2017digital]
Since the inception of DTs, research has focused on understanding the concept, the development of DT’s applications, or exploration of different implementation technologies used. An example is the work of Liu et al. [DT_concept_tech_applications] that analyses the concept of a DT, the technologies used, and DTs’ main industrial applications. That research is based on systematic literature review (SLR) analysing over 200 publications.
Tao et al. [Tao_DT-in-Industry] analyses the state-of-the-art in development and applications of DTs, aiming to outline key technologies enabling DT development, classify current and future applications, and lay out possible gaps and challenges. The research used an SLR, analyzing 50 papers and eight patents.
Sharma et al. [DT_state_art_1], based on an SLR of over 80 papers, analysed and proposed solutions for the research and implementation gaps of DT technology, such as IoT (internet of things), machine learning and data. This research concluded that regulation and security mechanisms are essential for the proper implementation of DTs due to its cross-domain nature. They also concluded that there are multiple technical and domain specific challenges that require more research to be resolved.
Gürdür et al. [interviews_DT_Infrastructure] explores how DTs can help the infrastructure industry. The research methodology used semi-structured expert interviews with non-technical executives from industry in the UK. This approach allowed the researchers to collect their opinions, related to non-technical challenges, on the value of DTs.
Dalibor et al. [Cross-Domain_SystMapping] conducted an SLR on 356 papers, analysing DTs with a bottom-up approach exploring different implementations to investigate expected DT properties and how DTs are deployed, operated and evaluated. In addition, the authors developed a DT feature model. They explored different implementation techniques, tooling and development processes.
@@ 3 Motivation and contribution
Although Grieves [grieves2017digital] introduced the concept of DT more than two decades ago, practitioners are discovering its potential only recently, resulting into an increase of activities and practices around DTs. An increasing number of DTs are being developed and used in domains such as automotive, healthcare, manufacturing, and construction. The application of these DTs extends across a wide spectrum of use-cases, ranging from basic monitoring to highly intricate control systems and predictive maintenance applications. Due to the diversity of use-cases and domains involved, the tools, methods, and practices around the design, development, and maintenance of DTs vary significantly. The central objective of this research is to understand and study this diversity. Our work concentrates on the design, development, and maintenance of the virtual part of a DT, which we refer to as virtual entity (VE) earlier.
VE, which is often described as digital replication of certain capabilities of a physical entity, is essentially a complex software- and data-intensive system that encompasses various heterogeneous and multi-domain models [Wright2020, Tao_DT-in-Industry]. To maintain the continued operation of a DT, the corresponding models must be kept consistent with each other, they need to be orchestrated properly, and finally, the dynamic behaviour of both the individual models and the DT as a whole must be validated. Although these tasks are well-known and researched within the broader model-driven engineering (MDE) community, we are unaware of any work focusing on these topics in the context of DTs. This deficiency motivates our work to interview 19 DT practitioners from industry and academia, in order to acquire insight into current challenges, methods and practices related to the three previously mentioned activities (i.e., KC1–3) in the context of DTs.
@@ 4 Research methodology
Considering the exploratory nature of our research questions we opted for semi-structured interviews [magaldi2020semi]. This provided sufficient flexibility for the participants to express themselves while allowing us to collect data on our topics of interest. In this section, we introduce and explain our research methodology, following Strandberg’s interview lifecycle [Strandberg2019EthicalEngineering], with the steps depicted in Figure 3. We expand on each step and explain the related activities in the following subsections.
@@ 4.1.2 Questionnaire design
We prepared a questionnaire to act as a guideline to keep the discussion in our semi-structured interviews focused. It was developed in line with the Interview Protocol Refinement (IPR) framework [castillo2016preparing], comprising four phases:
@@
    Constructing an inquiry-based conversation: We categorized the interview questions into (1) background, (2) key, and (3) concluding ones. Based on this categorization and suggestions of Hove and Anda [Hove2005ExperiencesResearch], we sorted them and rephrased some, enabling an inquiry-based conversation.
@@
Table 1 shows the resulting set of interview questions and their associations with the RQs. All the interview questions except for the first two are connected to one or more RQs. These two questions allowed us to start the conversation, get acquainted with the interviewee, and contributed to a conversation-like interview. Furthermore, with the final question we asked the interviewees’ opinion on Tao et al.’s 5-D DT model [tao2017digital] (see Section 2). While asking this question, we showed an image of the 5-D model and briefly explained it. To avoid influencing the interviewees during the rest of the interview, we intentionally asked this question at the end.
@@
Research questions (RQs) $\rightarrow$ 1 2 3 4 5 6 7 8
@@
1. What is the domain of your company and what kind of services does it provide?
2. How long have you been working and what is your current role?
@@
3. What is your understanding of a DT? $\times$
4. How are you involved in the development or usage of the DT?
5. What problems are you solving with your DT? $\times$
6. Is your DT for the entire process/system or a specific part? $\times$
7. What are the main parts of your DT? Could you shortly describe their role? $\times$ $\times$
8. Is there a physical counterpart of your DT? Does it communicate with the digital world? If yes, how? $\times$ $\times$
9. Did you build your DT from scratch or you reused some of the things that already existed? What kind of issues did you face if you reused something or by building from scratch? $\times$ $\times$
10. How is the data exchange between the models (and physical system) specified? $\times$ $\times$ $\times$
11. How is the sequence of execution of the models in your DT? $\times$
12. What issues, if any, did you face with the overall collective execution of models in your DT? $\times$
13. Which platforms/tools are you using to develop the digital twin? $\times$ $\times$ $\times$ $\times$
14. How often is the DT updated for bug fixes, improvement or similar reasons? $\times$
15. After each update did you face integration issues? If yes, what kind? $\times$ $\times$
16. How do you ensure various software elements can work together specially if multiple tools were used for development? $\times$ $\times$ $\times$ $\times$ $\times$
17. What properties/parameters did you validate to ensure an overall consistent behavior of your DT? $\times$
18. What tools and techniques did you use to validate these properties and parameters? $\times$ $\times$
19. What do you consider to be the general characteristics of your DT? $\times$
@@
20. How do you see the DT evolve in the future to solve additional problems? $\times$
21. What is your opinion on the 5-dimensional DT model from Tao et al. [tao2017digital]? $\times$
@@ 4.5 Data analysis
We further analysed the transcripts based on the thematic qualitative analysis methodology [Kiger2020], the activities related to which are explained in this section. Each of these activities utilize outcomes from one or more of the previous activities. Figure 4 depicts the relationships among the activities and the related outcomes.
We used LaMa [lama2023], a web-based tool for collaborative labeling and thematic analysis, to collaborate on this analysis. To restrict access and ensure data privacy, we deployed this platform locally.
@@ 4.5.2 Labeling of artifact and topic generation
After generating the artifacts for all transcripts, we labeled them using LaMa [lama2023]. We define a label as a short text that sufficiently captures the core message of an artifact. To reduce bias in labeling, each artifact was labeled by two labelers. We resolved conflicting labels by agreeing on one label through discussion. During the labeling process, the labeler could use an existing label or create a new one. In LaMa, these labels were accompanied by a description explaining how and when a label should be used, which was crucial for reuse of existing labels. Moreover, while labeling we encountered artifacts that lacked sufficient information or context. We labeled these artifacts with two predefined labels: No value or Not understandable. An example is an artefact discussing an interviewee research policy in relation with their clients. This artefact was classified as No value, since it does not discuss any information related to DT development.
@@ 4.6 Replication package
The replication package for this interview study is available at [muctadir_2023_10187933]. Due to the agreement made with our interviewees, we must keep the interview transcripts confidential. Although this significantly constraints on the reproducibility and validation of the findings presented in this paper, we believe that the provided package can still significantly contribute to similar research endeavours. This package includes:
@@ 6.1.1 Virtual representation of an entity
Table 3: Correlation between DT components from interviewees and 5D model [tao2017digital] addressed by industrial (#I) and academic (#A) interviewees.
@@ 6.1.2 Components of a DT
  • Purpose: Thirteen interviewees expressed that a DT should have a purpose and some mentioned that this purpose is the driving factor for DTs to be developed. On the contrary, one industrial interviewee explicitly mentioned that DTs should not have a specific purpose. This interviewee further discussed that DTs should not be developed with a specific purpose and they have a purposeless existence, which is in stark contrast to what the thirteen interviewees mentioned above. This interviewee clarified that once the DT is developed, it can be used for whichever purpose is needed. The purpose mentioned by the different interviewees can be correlated to the ‘services’ component of Tao et al. [tao2017digital]’s 5D model of DTs (see Section 2).
@@ 6.1.3 Relation to the 5D DT model
As explained in Section 4.1.2, we collected and analyzed the opinions of our interviewees on the 5D DT model by Tao et al. [tao2017digital] (explained in Section 2). This 5D DT model was selected in this research because it is a widely used model for DT representation and has been adopted for discussions on DTs by many. So, we wanted to understand the opinions of interviewees on this highly used and adopted model. In this section, we present our observations.
@@
Table 4: Agreement on 5D model of DT [tao2017digital] by interviewees from academia (#A) and industry (#I).